OpenStudy (anonymous):

Help please! A box in a college bookstore contains books, and each book in the box is a history book, an english book or science book. If one third of these books are history books and one sixth are english books, what fraction of the books are science books? A. 1/3 B. 1/2 C 2/3 D 3/4

OpenStudy (anonymous):

@wolf1728 help please!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

me either

OpenStudy (anonymous):

\[\frac{ 6 }{ 6 } = \frac{ 2 }{ 6 } - \frac{ 1 }{ 6 } - x\] solving for x gives you 1/2

OpenStudy (anonymous):

but how do you come with 1/6 or 2/6 if it only mentions 1/3 and 1/6

Directrix (directrix):

1/3 + 1/6 + S = 1 where 1/3 of the box is History, 1/6 is English and S is the fractional portion of Science books in the book. The number 1 represent the entire box. Get a common denominator of 6 for 1/3. 1/3 = ?/6 That would be 2 or 2/6 1/3 + 1/6 + S = 1 2/6 + 1/6 + S = 1 3/6 + S = 1 1/2 + S = 1 Note that 3/6 = 1/2 S = 1 - 1/2 S = ? fraction of the box designated as Science books. @Kfsv10